Real estate ads suggest that 58 % of homes for sale have garages,
16 % have swimming pools, and
9 % have both features.
What is the probability that a home for sale has
a) a pool or a garage?
Answer =  %
b) neither a pool nor a garage?
Answer =  %
c) a pool but no garage?
Answer =
Answer #1

a) P(A pool or a garage) = 0.58 + 0.16 - 0.09 = 0.65

b) P(Neither) = 1 - P(A pool or a garage) = 1 - 0.65 = 0.35

c) P(A pool but no garage) = P(Pool) - P(Both) = 0.16 - 0.09 = 0.07